What Is 3 Divided By 1 1/3
Found 2 solutions by checkley79, lenny460:
Answer by checkley79(3050)

3/1.333=2.25 or 2 1/4 ans.

Answer by lenny460(766)

3 divide by 1 1/3
3 / 4/3
3 * 3/4
3 * 3 = 9
9/4 = 2 1/4 or 2.25

The Answer:

2 1/4 or 2.25


obuong3@aol.com